# Causal Effect Estimation under Networked Interference without Networked Unconfoundedness Assumption

### v1

@@ 6 Experiments
In this section, we validate the proposed method on two commonly used semisynthetic datasets. In detail, we verify the effectiveness of our algorithm and further evaluate the correctness of the analysis with the help of semisynthetic datasets.
@@
Table 1: Experimental results on Flickr(homo) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
@@
Ours_w/o_IPM $0.0359_{\pm 0.0262}$ $\textbf{0.0133}_{\pm 0.0050}$ $0.0598_{\pm 0.0366}$ $0.0394_{\pm 0.0262}$ $\textbf{0.0123}_{\pm 0.0060}$ $0.0602_{\pm 0.0355}$ $0.0410_{\pm 0.0228}$ $\textbf{0.0155}_{\pm 0.0033}$ $0.0643_{\pm 0.0337}$ $0.0424_{\pm 0.0242}$ $\textbf{0.0142}_{\pm 0.0044}$ $0.0632_{\pm 0.0334}$
Ours $\textbf{0.0296}_{\pm 0.0219}$ $0.0252_{\pm 0.0208}$ $\textbf{0.0266}_{\pm 0.0208}$ $\textbf{0.0289}_{\pm 0.0208}$ $0.0252_{\pm 0.0208}$ $\textbf{0.0260}_{\pm 0.0200}$ $\textbf{0.0297}_{\pm 0.0220}$ $0.0252_{\pm 0.0208}$ $\textbf{0.0267}_{\pm 0.0209}$ $\textbf{0.0289}_{\pm 0.0209}$ $0.0252_{\pm 0.0208}$ $\textbf{0.0261}_{\pm 0.0201}$
6.1 Experimental Setup
Table 2: Experimental results on Flickr(hete) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
@@ 6.1 Experimental Setup
TNet $0.0411_{\pm 0.0238}$ $0.0206_{\pm 0.0073}$ $0.0282_{\pm 0.0297}$ $0.0417_{\pm 0.0237}$ $0.0196_{\pm 0.0098}$ $0.0268_{\pm 0.0314}$ $0.0936_{\pm 0.0170}$ $0.0338_{\pm 0.0065}$ $0.1360_{\pm 0.0210}$ $0.0950_{\pm 0.0157}$ $0.0361_{\pm 0.0074}$ $0.1415_{\pm 0.0223}$
Ours_w/o_IPM $\textbf{0.0296}_{\pm 0.0248}$ $\textbf{0.0125}_{\pm 0.0131}$ $0.0581_{\pm 0.0321}$ $\textbf{0.0254}_{\pm 0.0212}$ $\textbf{0.0171}_{\pm 0.0195}$ $0.0437_{\pm 0.0241}$ $0.0929_{\pm 0.0181}$ $0.0361_{\pm 0.0114}$ $0.1503_{\pm 0.0221}$ $\textbf{0.0910}_{\pm 0.0153}$ $0.0393_{\pm 0.0149}$ $0.1480_{\pm 0.0172}$
Ours $0.0377_{\pm 0.0207}$ $0.0206_{\pm 0.0157}$ $\textbf{0.0228}_{\pm 0.0058}$ $0.0384_{\pm 0.0225}$ $0.0226_{\pm 0.0177}$ $\textbf{0.0197}_{\pm 0.0088}$ $\textbf{0.0922}_{\pm 0.0115}$ $\textbf{0.0353}_{\pm 0.0121}$ $\textbf{0.1326}_{\pm 0.0161}$ $0.0941_{\pm 0.0100}$ $\textbf{0.0390}_{\pm 0.0135}$ $\textbf{0.1379}_{\pm 0.0164}$
Datasets We consider two wildly used semisynthetic datasets BlogCatalog and Flickr to verify the effectiveness of our estimator. We further use a synthetic dataset to validate the correctness of our theories, i.e., whether our method can correctly recover latent confounders.
Following existing works (Jiang & Sun, 2022; Guo et al., 2020; Ma et al., 2021; Chen et al., 2024), we use two semisynthetic datasets to evaluate our proposed method:
  • Flickr is an online social network where users can share images and videos. In this dataset, each unit is a user and each edge is the social relationship between units. The features are the list of tags of units’ interests.
We reuse the original covariates as the latent confounders and then divide them into $u^{i},u^{c},u^{n}$. We generate the proxies $x$ using $x_{i}=w_{1}u_{i}+e_{i}$ where $w_{1}$ are randomly sampled from Uniform distribution $\mathcal{U}(0.5,1)$ and $e_{x,i}$ is standard Gaussian noise. Then given the latent confounder $u^{i}_{i},u^{c}_{i},u^{n}_{i}$ of unit $i$, the treatments are simulated by
$t_{i}=\begin{cases}1&\text{if \quad$tpt_{i}>\overline{tpt}$},\\ 0&\text{else},\end{cases}$
where $\overline{tpt}$ is the average of all $tpt_{i}$, and $tpt_{i}=pt_{i}+pt_{\mathcal{N}_{i}}$, and $pt_{i}=Sigmoid(w_{2}\times[u^{i}_{i},u^{c}_{i}])$, and $pt_{\mathcal{N}_{i}}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j}^{j\in\mathcal{N}_{i}}Sigmoid(w_{3}\times[u^{i}_{c},u^{n}_{i}])$ serves as the neighbors influences. Here $w_{2}$ and $w_{3}$ are randomly generated weight vectors that mimic the causal mechanism from the latent confounders to treatments. Then, $z_{i}$ can be directly obtained by network topology $E$ and $t_{\mathcal{N}_{i}}$.
We then modify the data generation of outcome $y$ in Jiang & Sun (2022):
$\small y_{i}(t_{i},z_{i})=t_{i}+z_{i}+po_{i}+0.5\times po_{\mathcal{N}_{i}}+e_{y,i},$
Also following Chen et al. (2024), we consider the data generation of outcome $y$ with heterogeneous effect:
$\small y_{i}$ $(t_{i},z_{i})=t_{i}+z_{i}+po_{i}+0.5\times po_{\mathcal{N}_{i}}$
$+t_{i}(po_{i}+0.5\times po_{\mathcal{N}_{i}})+t_{i}(0.5\times po_{i}+po_{\mathcal{N}_{i}})+e_{y,i},$
which is denoted as BC(hete) and Flickr(hete).
Due to the space limit, we leave the detailed data generation process of the synthetic dataset in Appendix C.2.
Baselines We denote our method as Ours and a variant without IPM loss as Ours_w/o_IPM. We compare our methods with several state-of-the-art baselines  [^2] The details are in Appendix C.1. Our code will be available upon acceptance. Following Chen et al. (2024), we modify TARNET, CFR (Johansson et al., 2021) and ND (Guo et al., 2020) by additionally inputting the exposure $z_{i}$, denoted as TARNET+z, CFR+z and ND+z respectively. We also consider several baselines that are designed for networked effect estimation under the same setting, including GEst (Ma & Tresp, 2021), NetEst (Jiang & Sun, 2022), and TNet (Chen et al., 2024).
Metrics In this paper, we use the Mean Absolute Error (MAE) on AME, ASE, and ATE as our metric, i.e., $\varepsilon_{average}=|\hat{\tau}-\tau|$, where $\tau$ and $\hat{\tau}$ are the average causal effect and estimated one. We also use the Rooted Precision in Estimation of Heterogeneous Effect on IME, ASE, and ITE, $\varepsilon_{individual}=\sqrt{\frac{1}{n}\Sigma_{i=1}^{n}(\hat{\tau}_{i}-\tau_{i})^{2}}$, where $\tau_{i}$ and $\hat{\tau}_{i}$ are the individual causal effect and estimated one. The mean and standard deviation of these metrics via $5$ times running are reported. Note that our main estimands are AME, ASE, and ATE in this paper.
6.2 Experimental Analyses
Figure 4: Visualization of recovered and ground-true latent confounders $U^{i},U^{c}$, and $U^{n}$.

### v2

@@ VI Experiments
In this section, we verify the effectiveness of our method, CaLaNet, and further evaluate the correctness of the theories using the commonly used semisynthetic datasets. In particular, we aim to answer the following research questions (RQs):
  • RQ1: How does the proposed method perform compared with existing methods?
  • RQ2: Can the proposed model correctly recover the latent confounders?
  • RQ3: Does our model stably perform well under different choices of hyperparameters?
@@
TABLE I: Experimental results on BC(homo) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
$\varepsilon_{average}$ $\varepsilon_{individual}$
Within Sample Out-of Sample Within Sample Out-of Sample
Methods AME ASE ATE AME ASE ATE IME ISE ITE IME ISE ITE
CFR+z $0.0788_{\pm 0.0096}$ $0.1157_{\pm 0.0076}$ $0.2323_{\pm 0.0106}$ $\underline{0.0770}_{\pm 0.0099}$ $0.1157_{\pm 0.0075}$ $0.2306_{\pm 0.0106}$ $\underline{0.0796}_{\pm 0.0091}$ $0.1158_{\pm 0.0076}$ $0.2325_{\pm 0.0106}$ $0.1000_{\pm 0.0388}$ $0.1157_{\pm 0.0075}$ $0.2405_{\pm 0.0166}$
ND+z $0.2375_{\pm 0.0450}$ $\underline{0.0316}_{\pm 0.0104}$ $0.0790_{\pm 0.0226}$ $0.2380_{\pm 0.0458}$ $\underline{0.0323}_{\pm 0.0122}$ $0.0768_{\pm 0.0254}$ $0.2377_{\pm 0.0448}$ $\underline{0.0321}_{\pm 0.0101}$ $0.0792_{\pm 0.0226}$ $0.2477_{\pm 0.0460}$ $\underline{0.0379}_{\pm 0.0099}$ $0.1068_{\pm 0.0172}$
NetEst $0.0989_{\pm 0.0735}$ ${0.0428}_{\pm 0.0487}$ $0.0596_{\pm 0.0291}$ $0.0995_{\pm 0.0725}$ $0.0425_{\pm 0.0487}$ $0.0599_{\pm 0.0286}$ $0.0992_{\pm 0.0732}$ $0.0431_{\pm 0.0485}$ $0.0602_{\pm 0.0290}$ $0.1130_{\pm 0.0617}$ $0.0553_{\pm 0.0443}$ $0.0660_{\pm 0.0254}$
RRNet $0.0884_{\pm 0.0495}$ $0.0445_{\pm 0.0158}$ $0.0768_{\pm 0.0239}$ $0.0892_{\pm 0.0505}$ $0.0447_{\pm 0.0160}$ $0.0782_{\pm 0.0253}$ $0.0915_{\pm 0.0462}$ $0.0452_{\pm 0.0153}$ $0.0865_{\pm 0.0260}$ $\underline{0.0917}_{\pm 0.0491}$ $0.0453_{\pm 0.0159}$ $0.0887_{\pm 0.0282}$
SPNet+z $\underline{0.0802}_{\pm 0.0632}$ $0.0446_{\pm 0.0155}$ $0.0570_{\pm 0.0188}$ $0.0931_{\pm 0.0607}$ $0.0394_{\pm 0.0215}$ $\underline{0.0385}_{\pm 0.0183}$ $0.1153_{\pm 0.0482}$ $0.0741_{\pm 0.0109}$ $0.1094_{\pm 0.0274}$ $0.1358_{\pm 0.0566}$ $0.0604_{\pm 0.0145}$ $0.0933_{\pm 0.0342}$
TNet $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $\underline{0.0473}_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $\underline{0.0473}_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $\underline{0.0473}_{\pm 0.0229}$
CaLaNet $\textbf{0.0551}_{\pm 0.0423}$ $\textbf{0.0279}_{\pm 0.0279}$ $\textbf{0.0288}_{\pm 0.0150}$ $\textbf{0.0549}_{\pm 0.0426}$ $\textbf{0.0278}_{\pm 0.0279}$ $\textbf{0.0288}_{\pm 0.0150}$ $\textbf{0.0554}_{\pm 0.0418}$ $\textbf{0.0279}_{\pm 0.0279}$ $\textbf{0.0289}_{\pm 0.0149}$ $\textbf{0.0551}_{\pm 0.0424}$ $\textbf{0.0278}_{\pm 0.0279}$ $\textbf{0.0288}_{\pm 0.0149}$
VI-A Experimental Setup
To begin with, we first introduce our experimental setup, including the used datasets, the compared baselines, and the used metrics.
VI-A1 Datasets
We consider two widely used semisynthetic datasets, BlogCatalog and Flickr, to verify the effectiveness of our estimator. We further use a synthetic dataset to validate the correctness of our theories, i.e., whether our method can correctly recover latent confounders.
Following existing works [25, 43, 47, 9], we use two semisynthetic datasets to evaluate our proposed method:
  • Flickr is an online social network where users can share images and videos. In this dataset, each unit is a user, and each edge is the social relationship between units. The features are the list of tags of units’ interests.
We reuse the original covariates as the latent confounders and then divide them into $u^{i},u^{c},u^{n}$. We generate the proxies $x_{i}$ using $x_{i}=w_{1}u_{i}+e_{i}$ where $w_{1}$ are randomly sampled from Uniform distribution $\mathcal{U}(0.5,1)$ and $e_{x,i}$ is standard Gaussian noise. Then given the latent confounder $u^{i}_{i},u^{c}_{i},u^{n}_{i}$ of unit $i$, the treatments are simulated by
[…]
where $\overline{tpt}$ is the average of all $tpt_{i}$, and $tpt_{i}=pt_{i}+pt_{\mathcal{N}_{i}}$, and $pt_{i}=Sigmoid(w_{2}\times[u^{i}_{i},u^{c}_{i}])$, and $pt_{\mathcal{N}_{i}}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j}^{j\in\mathcal{N}_{i}}Sigmoid(w_{3}\times[u^{i}_{c},u^{n}_{i}])$ serves as the neighbors influences. Here, $w_{2}$ and $w_{3}$ are randomly generated weight vectors that mimic the causal mechanism from the latent confounders to treatments. Then, $z_{i}$ can be directly obtained by network topology $E$ and $t_{\mathcal{N}_{i}}$.
We then modify the data generation of outcome $y$ in [25] using the latent confounders $u_{i}$ instead of observed $x_{i}$:
[…]
Also following [9], we consider the data generation of outcome $y$ with heterogeneous effects:
[…]
and the datasets are denoted as BC(hete) and Flickr(hete).
Due to the space limit, we leave the detailed data generation process of the synthetic dataset in Appendix E.
TABLE II: Experimental results on BC(hete) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
$\varepsilon_{average}$ $\varepsilon_{individual}$
Within Sample Out-of Sample Within Sample Out-of Sample
Methods AME ASE ATE AME ASE ATE IME ISE ITE IME ISE ITE
ND+z $0.4124_{\pm 0.0702}$ $\underline{0.0451}_{\pm 0.0201}$ $0.1330_{\pm 0.0205}$ $0.4111_{\pm 0.0737}$ $0.0486_{\pm 0.0206}$ $0.1326_{\pm 0.0261}$ $0.4226_{\pm 0.0673}$ $\underline{0.0562}_{\pm 0.0146}$ $0.1941_{\pm 0.0246}$ $0.4330_{\pm 0.0662}$ $0.0666_{\pm 0.0137}$ $0.2211_{\pm 0.0321}$
NetEst $0.2301_{\pm 0.0784}$ $0.0799_{\pm 0.0453}$ $0.1296_{\pm 0.0594}$ $0.2252_{\pm 0.0729}$ $0.0785_{\pm 0.0441}$ $0.1221_{\pm 0.0594}$ $0.2479_{\pm 0.0737}$ $0.0859_{\pm 0.0433}$ $0.1960_{\pm 0.0444}$ $0.2485_{\pm 0.0761}$ $0.0923_{\pm 0.0433}$ $0.1988_{\pm 0.0435}$
RRNet $0.1628_{\pm 0.1183}$ ${0.0476}_{\pm 0.0416}$ $0.0919_{\pm 0.0581}$ $0.1636_{\pm 0.1162}$ $\underline{0.0474}_{\pm 0.0424}$ $0.0961_{\pm 0.0640}$ $0.1967_{\pm 0.0979}$ $0.0591_{\pm 0.0364}$ $0.1719_{\pm 0.0485}$ $0.1974_{\pm 0.0960}$ $\underline{0.0595}_{\pm 0.0365}$ $0.1764_{\pm 0.0505}$
SPNet+z $\underline{0.0706}_{\pm 0.0738}$ $0.0714_{\pm 0.0256}$ $0.1169_{\pm 0.0317}$ $\underline{0.0805}_{\pm 0.0716}$ $0.0671_{\pm 0.0230}$ $0.0841_{\pm 0.0419}$ $0.1743_{\pm 0.0338}$ $0.1054_{\pm 0.0138}$ $0.2209_{\pm 0.0288}$ $\underline{0.1901}_{\pm 0.0480}$ $0.0906_{\pm 0.0190}$ $0.2012_{\pm 0.0480}$
TNet $0.1216_{\pm 0.0864}$ $0.0537_{\pm 0.0524}$ $\underline{0.0429}_{\pm 0.0301}$ $0.1257_{\pm 0.0727}$ $0.0537_{\pm 0.0511}$ $0.0481_{\pm 0.0269}$ $\underline{0.1731}_{\pm 0.0450}$ $0.0655_{\pm 0.0465}$ $\textbf{0.1458}_{\pm 0.0175}$ $0.1915_{\pm 0.0542}$ $0.0650_{\pm 0.0458}$ $\underline{0.1740}_{\pm 0.0621}$
CaLaNet $\textbf{0.0701}_{\pm 0.0571}$ $\textbf{0.0321}_{\pm 0.0239}$ $\textbf{0.0401}_{\pm 0.0264}$ $\textbf{0.0691}_{\pm 0.0550}$ $\textbf{0.0305}_{\pm 0.0243}$ $\textbf{0.0396}_{\pm 0.0265}$ $\textbf{0.1224}_{\pm 0.0310}$ $\textbf{0.0458}_{\pm 0.0181}$ $\underline{0.1481}_{\pm 0.0210}$ $\textbf{0.1216}_{\pm 0.0300}$ $\textbf{0.0452}_{\pm 0.0178}$ $\textbf{0.1493}_{\pm 0.0185}$
TABLE III: Experimental results on Flickr(homo) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
@@ VI-A1 Datasets
RRNet $\underline{0.0296}_{\pm 0.0123}$ ${0.0251}_{\pm 0.0172}$ $\underline{0.0199}_{\pm 0.0179}$ $\underline{0.0296}_{\pm 0.0123}$ ${0.0251}_{\pm 0.0172}$ $\underline{0.0199}_{\pm 0.0179}$ $\underline{0.0296}_{\pm 0.0123}$ ${0.0251}_{\pm 0.0172}$ $\underline{0.0199}_{\pm 0.0179}$ $\underline{0.0296}_{\pm 0.0123}$ $\underline{0.0251}_{\pm 0.0172}$ $\underline{0.0199}_{\pm 0.0179}$
SPNet +z $0.0366_{\pm 0.0359}$ $0.0604_{\pm 0.0371}$ $0.1267_{\pm 0.0655}$ $0.0451_{\pm 0.0332}$ $0.0445_{\pm 0.0300}$ $0.0865_{\pm 0.0410}$ $0.0639_{\pm 0.0273}$ $0.0980_{\pm 0.0269}$ $0.1877_{\pm 0.0546}$ $0.0834_{\pm 0.0342}$ $0.0694_{\pm 0.0234}$ $0.1309_{\pm 0.0500}$
@@
CaLaNet $\textbf{0.0223}_{\pm 0.0159}$ $\textbf{0.0147}_{\pm 0.0068}$ $\textbf{0.0151}_{\pm 0.0127}$ $\textbf{0.0222}_{\pm 0.0158}$ $\textbf{0.0148}_{\pm 0.0069}$ $\textbf{0.0154}_{\pm 0.0128}$ $\textbf{0.0223}_{\pm 0.0159}$ $\textbf{0.0149}_{\pm 0.0069}$ $\textbf{0.0152}_{\pm 0.0127}$ $\textbf{0.0223}_{\pm 0.0158}$ $\textbf{0.0148}_{\pm 0.0068}$ $\textbf{0.0155}_{\pm 0.0129}$
VI-A2 Baselines
We denote our method as CaLaNet  [^2] Our code will be available upon acceptance.. We compare CaLaNet with several state-of-the-art baselines as follows:
  • TARNET+z: Original Tarnet [48] uses two-heads neural networks, serving as T-Learner-like estimator, to estimate causal effects under no interference assumption. We modify TARNET by additionally inputting the exposure $z_{i}$.
  • CFR+z: Original CFR [48] uses two-heads neural networks with an MMD term to achieve counterfactual regression under no interference assumption. We modify CFR by additionally inputting the exposure $z_{i}$.
  • ND+z: Original ND [43] propose network deconfounder framework by using network information under no interference assumption. We modify ND by additionally inputting the exposure $z_{i}$.
  • GEst [6]: GEst, based on CFR, uses GCN to aggregate the features of neighbors and input the exposure $z_{i}$ to estimate causal effects under networked interference.
  • NetEst [25]: NetEst learns balanced representation via adversarial learning for networked causal effect estimation.
  • RRNet [7]: RRNet combines the representation learning and reweighting techniques in its network to estimate causal effects under interference.
  • SPNet+z [29]: SPNet estimates the effect of the individual treatment $t_{i}$while keeping the neighbor exposure $z_{i}$ unchanged. We adapt SPNet to our setting by incorporating $z_{i}$ as an additional input.
  • TNet [9]: TNet utilizes targeted learning techniques into its neural networks model to estimate causal effects under networked interference in a double robust manner.
VI-A3 Metrics
In this paper, we use the Mean Absolute Error (MAE) on AME, ASE, and ATE as our metric, i.e., $\varepsilon_{average}=|\hat{\tau}-\tau|$, where $\tau$ and $\hat{\tau}$ are the average causal effect and estimated one. We also use the Rooted Precision in Estimation of Heterogeneous Effect on IME, ASE, and ITE, $\varepsilon_{individual}=\sqrt{\frac{1}{n}\Sigma_{i=1}^{n}(\hat{\tau}_{i}-\tau_{i})^{2}}$, where $\tau_{i}$ and $\hat{\tau}_{i}$ are the individual causal effect and estimated one. The mean and standard deviation of these metrics via 555 times running are reported. Note that our main estimands are AME, ASE, and ATE in this paper.
TABLE IV: Experimental results on Flickr(hete) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
@@ VI-A3 Metrics
RRNET $0.0441_{\pm 0.0196}$ ${0.0280}_{\pm 0.0153}$ $\underline{0.0367}_{\pm 0.0243}$ $0.0446_{\pm 0.0183}$ $0.0292_{\pm 0.0165}$ ${0.0329}_{\pm 0.0255}$ $0.0940_{\pm 0.0166}$ $0.0401_{\pm 0.0105}$ $0.1373_{\pm 0.0188}$ $0.0952_{\pm 0.0153}$ $0.0436_{\pm 0.0100}$ $0.1419_{\pm 0.0202}$
SPNet+z $0.0429_{\pm 0.0461}$ $0.0885_{\pm 0.0509}$ $0.2359_{\pm 0.0724}$ $0.0509_{\pm 0.0404}$ $0.0555_{\pm 0.0264}$ $0.1483_{\pm 0.0573}$ $0.1296_{\pm 0.0241}$ $0.1593_{\pm 0.0314}$ $0.3538_{\pm 0.0513}$ $0.1827_{\pm 0.1051}$ $0.1053_{\pm 0.0208}$ $0.2879_{\pm 0.0823}$
TNet $\underline{0.0411}_{\pm 0.0238}$ $\underline{0.0206}_{\pm 0.0073}$ $\underline{0.0282}_{\pm 0.0297}$ $\underline{0.0417}_{\pm 0.0237}$ $\underline{0.0196}_{\pm 0.0098}$ $\underline{0.0268}_{\pm 0.0314}$ $\underline{0.0936}_{\pm 0.0170}$ $\underline{0.0338}_{\pm 0.0065}$ $\underline{0.1360}_{\pm 0.0210}$ $\underline{0.0950}_{\pm 0.0157}$ $\underline{0.0361}_{\pm 0.0074}$ $\underline{0.1415}_{\pm 0.0223}$
CaLaNet $\textbf{0.0283}_{\pm 0.0223}$ $\textbf{0.0135}_{\pm 0.0070}$ $\textbf{0.0257}_{\pm 0.0099}$ $\textbf{0.0316}_{\pm 0.0206}$ $\textbf{0.0146}_{\pm 0.0074}$ $\textbf{0.0217}_{\pm 0.0091}$ $\textbf{0.0892}_{\pm 0.0111}$ $\textbf{0.0302}_{\pm 0.0053}$ $\textbf{0.1334}_{\pm 0.0144}$ $\textbf{0.0911}_{\pm 0.0099}$ $\textbf{0.0334}_{\pm 0.0055}$ $\textbf{0.1383}_{\pm 0.0151}$
VI-B Experimental Analyses
VI-B1 RQ1: Effectiveness of CaLaNet